# Doc2SAR: A Synergistic Framework for High-Fidelity Extraction of Structure-Activity Relationships from Scientific Documents

### v1

@@ Appendix E Web Application
Furthermore, recognizing that automated systems may occasionally require refinement, the web application incorporates an interactive editing interface (Figure 10(c)). Users can directly modify extracted coreferences, SMILES strings, or activity values. For SMILES correction, the interface supports direct text editing, drawing the chemical structure using an embedded molecular editor, or even uploading a corrected molecular image from which the SMILES can be re-derived. Finally, the curated and validated extraction results can be conveniently downloaded by the user in CSV format for downstream analysis and integration into their research workflows (Figure 10(d)). A live demonstration of the web application is accessible at https://uni-finder.dp.tech/molecular-extract.

### v2

@@ Appendix E Web Application
Furthermore, recognizing that automated systems may occasionally require refinement, the web application incorporates an interactive editing interface (Figure 10(c)). Users can directly modify extracted coreferences, SMILES strings, or activity values. For SMILES correction, the interface supports direct text editing, drawing the chemical structure using an embedded molecular editor, or even uploading a corrected molecular image from which the SMILES can be re-derived. Finally, the curated and validated extraction results can be conveniently downloaded by the user in CSV format for downstream analysis and integration into their research workflows (Figure 10(d)).
@@
Table 4: DocSAR-200 Data List
Literature Patent
10.1021_acsmedchemlett.4c00408 10.1021_acsmedchemlett.3c00180 10.1016_j.ejmech.2015.11.043 BR112016001107B1 US20160355824A1 CN102482273B
10.1021_acsmedchemlett.3c00475 10.1021_jm025553u 10.1021_acsmedchemlett.4c00133 US11155555 CA3012408C ES2988680T3
10.1016_j.bmcl.2017.03.030 10.1021_acsmedchemlett.3c00222 10.1021_acsmedchemlett.4c00261 CA3114414A1 AU2015238300C1 US10100308
10.1021_acsmedchemlett.4c00230 10.1021_acsmedchemlett.4c00168 10.1021_jm050499d CN102089304A CN103402511B US20110046151A1
10.1021_acsmedchemlett.3c00544 10.1016_j.ejmech.2023.116113 10.1021_jm049503w JP2018519304A CA2982708C US10435414
10.1021_acsmedchemlett.4c00195 10.1021_jm0600139 10.1021_acsmedchemlett.4c00286 CN102159555A DK3122742T3 US20130172386A1
10.1021_acsmedchemlett.3c00444 10.1021_acsmedchemlett.3c00507 10.1021_jm800936s ES2675791T3 WO2014113407A2 CN102574816A
10.1016_j.ejmech.2023.116098 10.1021_jm990622z 10.1021_acsmedchemlett.3c00312 TW201408641A WO2008134474A2 KR20230154230A
10.1021_acsmedchemlett.4c00331 10.1021_acsmedchemlett.4c00390 10.1021_acsmedchemlett.4c00329 ES2904513T3 US20170298047A1 BRPI0909017A2
10.1021_jm049660v 10.1021_acsmedchemlett.3c00392 10.1021_jm0155042 US10435407 EA040922B1 CN105793260B
10.1021_jm970540f 10.1021_acsmedchemlett.4c00076 10.1002_cmdc.202400302 EP3408265B1 DK2638031T3 US20230303530A1
10.1021_jm050413g 10.1021_jm701228e 10.1021_acsmedchemlett.2c00406 AU2022263269A1 BR112016021507B1 CA2479109C
10.1021_jm801101z 10.1021_acsmedchemlett.3c00506 10.1016_j.bmcl.2017.07.029 KR102306071B1 EP3464318B1 CA2600052A1
10.1021_jm015573g 10.1021_acsmedchemlett.3c00308 10.1021_jm060484v US8598357 AU2017216212B2 US8703811
10.1021_acsmedchemlett.3c00494 10.1021_jm990577v 10.1021_acsmedchemlett.3c00249 ES2357988T3 US20220056009A1 CN1659133A
10.1021_acsmedchemlett.4c00046 10.1021_acsmedchemlett.4c00177 10.1021_acsmedchemlett.4c00056 EP2094694B1 US20190209539A1 US20240092754A1
10.1016_j.bmcl.2024.129931 10.1021_acsmedchemlett.4c00217 10.1021_acsmedchemlett.4c00137 TW201024308A EP3271333B1 ES2577829T3
10.1021_acsmedchemlett.3c00393 10.1021_acsmedchemlett.4c00285 10.1021_acsmedchemlett.4c00414 ES2626908T3 AU2022230795A1 JP2024540522A
10.1021_acsmedchemlett.4c00119 10.1016_j.bmcl.2024.130012 10.1021_acsmedchemlett.3c00236 EA031639B1 US20230192660A1 CN109153640B
10.1021_acsmedchemlett.3c00234 10.1021_acsmedchemlett.4c00303 10.1021_jm0002679 CN103080089A CN107531683B AU2017281903B2
10.1016_j.bmcl.2024.129725 10.1021_jm049156q 10.1021_acsmedchemlett.4c00356 US11883500 US9006271 US10406157
10.1016_j.bmcl.2005.05.013 10.1021_acsmedchemlett.4c00332 10.1021_acsmedchemlett.3c00564 AU2012217150A1 EP2256117A1 AU2007321019B2
10.1016_j.bmcl.2017.06.049 10.1021_acsmedchemlett.3c00389 10.1002_cmdc.200900097 JP6898933B2 ES2991765T3 US20170209432A1
10.1021_acsmedchemlett.3c00323 10.1021_acsmedchemlett.4c00001 10.1021_acsmedchemlett.3c00177 EA035188B1 CN108434147A CN107849044A
10.1021_acsmedchemlett.4c00118 10.1021_acsmedchemlett.5c00128 10.1021_acsmedchemlett.3c00441 EP3552630A1 CA3238252A1 US20200255401A1
10.1021_acsmedchemlett.3c00271 10.1021_acsmedchemlett.5c00107 10.1016_s0223-5234(00)00142-2 US20220056540A1 US20190308987A1 US9132145
10.1021_acsmedchemlett.4c00304 10.1021_acsmedchemlett.3c00324 10.1016_j.bmcl.2024.129965 EP1716145B1 US20180008587A1 ES2546430T3
10.1021_acsmedchemlett.3c00235 10.1021_acsmedchemlett.2c00395 10.1016_j.bmcl.2012.04.036 EP2091939B1 EA046921B1 US20190071444A1
10.1021_jm0256068 10.1021_acsmedchemlett.4c00228 10.1021_acsmedchemlett.4c00343 JP2013530954A CN108290866A CA2806647C
10.1021_acsmedchemlett.3c00554 10.1021_jm980053f 10.1021_jm0610245 CA2767372A1 AU2016247476B2 US20180208570A1
10.1021_acsmedchemlett.3c00477 10.1016_j.bmcl.2018.07.049 10.1021_acsmedchemlett.4c00011 US20210380533A1 AU2003212335B8 JP6990657B2
10.1021_acsmedchemlett.4c00262 10.1021_acsmedchemlett.3c00474 10.1021_acsmedchemlett.3c00458 JP7002335B2 EP3455221B1 US20230203011A1
10.1002_cmdc.202100398 10.1021_acsmedchemlett.3c00473 10.1016_j.bmcl.2025.130230